# A -ray-emitting blazar B3 1239+376 at = 3.82 identified in a multi-wavelength context

### v1

@@ III Discussions and Summary
The simple leptonic model can reasonably reproduce the both SEDs, see Figure 7. The corresponding input parameters are listed in Table 1. The radio data points are not considered in the jet modeling because they are likely from an extended region rather than the compact blob, and the radio emissions are seriously suffered by the SSA. Our jet modeling analyses suggest that B3 1239+376 is likely a low-synchrotron-peaked source (i.e., Hz, Padovani & Giommi 1995; Abdo etĀ al. 2010). In the high flux state, the synchrotron bump rises and shifts close to the infrared domain, resulting the increase of the flux densities there. On the other hand, the high energy bump moves into the detection energy range of Fermi-LAT at the same time, which leads to the emergence of the -ray signal. Theoretically, the injection of fresh emitting electrons in the dissipation region, as well as the enhancement of the magnetic field intensity, could account for the broadband variations. Therefore, it is not surprising that input parameters such as and B for the high flux state SED are larger than that of the other SED. The Compton dominance of B3 1239+376 is indicated as 10, consistent with that found in other high- blazars (e.g., Ghisellini etĀ al., 2011; Ackermann etĀ al., 2017). Interestingly, even in the low flux state, the X-ray emission of B3 1239+376 is luminous, which is also seen in several -ray quiet high- blazars (Marcotulli etĀ al., 2020, 2025). Together with the hard X-ray spectra, blazars are believed as powerful MeV beacons in the early cosmos. Meanwhile, an important jet parameter, the bulk Lorentz factor (), is worthy to be mentioned specially. In the classic orientation based AGN unified scheme, the number ratio of the jetted AGNs between the aligned and the apparent populations is suggested as . Thus, is a key parameter to determine the number density of jetted AGNs in the early cosmic time. The value for B3 1239+376 obtained here is 15, in agreement with the results (i. e., 1020) provided for other similar sources (e.g., Ghisellini etĀ al., 2011; Ackermann etĀ al., 2017; Liao etĀ al., 2020). In addition to the jet SED modeling studies, other approaches are used to probe the jet Lorentz factor of high- sources. Searching for the potential evidence of fast variability, especially in rays, can put a severe constraint on the value (e.g., Li etĀ al., 2018). Moreover, detections of the proper motion of the radio jets provide a direct estimation of the essential properties of the jet (e.g., Zhang etĀ al., 2020; Krezinger etĀ al., 2024). Considering the recent -ray activity of B3 1239+376 revealed by Fermi-LAT, quick follow-up radio interferometry observations that would reveal its jet kinematics are strongly encouraged.
@@
In summary, based on detailed multi-wavelength investigations, B3 1239+376 ( = 3.82) is identified as a -ray emitter, and hence it is concluded as the third most distant -ray-emitting blazar so far, underneath B3 1428+422 ( = 4.72) and GB 1508+5714 ( = 4.3). Analysis of Fermi-LAT data yields a significantly detected (7.7 globally) -ray source cospatial with B3 1239+376 in a half-year period of 2025. It is known as the unique blazar (candidate) within the location uncertainty and an association probability of 0.91 is estimated, which suggest a physical connection between the -ray source and B3 1239+376. Meanwhile, indications of prior -ray activities are also found. Although for the two epochs each alone the statistics are limited, a joint analysis returns a boosted detection significance and the B3 1239+376 remains within the -ray localization error radius. Moreover, in a multi-wavelength perspective, the -ray and infrared fluxes contemporaneously brighten. In particular, emissions in these two windows reach up to the maximal level at the same time, under a total 16-year monitoring. The temporal coincidence helps us to determine the association relationship. Interestingly, accompanying radio core flux increase until Mar. 2018 is observed. In addition to the temporal behaviors, broadband SEDs of B3 1239+376 are plotted and described by a one-zone leptonic radiation scenario. Ejections of fresh emitting electrons into the dissipation region together with an increase of the magnetic field intensity may be responsible for the variations. Comparison between B3 1239+376 and other distant blazars are carried out. The soft -ray spectrum, luminous X-ray and -ray emissions, and variability across different bands, are in agreement with that found in other similar sources. To further investigate its jet properties, future multi-wavelength campaigns are urged.

### v2

@@ IV Discussions and Summary
Our jet modeling analyses suggest that B3 1239+376 is likely a low-synchrotron-peaked source (i.e., Hz, Padovani and Giommi 1995; Abdo et al. 2010). In the high flux state, the synchrotron bump rises and shifts close to the infrared domain, resulting the increase of the flux densities there. On the other hand, the high energy bump moves into the detection energy range of Fermi-LAT at the same time, which leads to the emergence of the -ray signal. Theoretically, the injection of fresh emitting electrons in the dissipation region, as well as the enhancement of the magnetic field intensity, could account for the broadband variations. Therefore, it is not surprising that input parameters such as and B for the high flux state SED are larger than that of the other SED. Recent multi-wavelength campaign studying on high- blazars beyond redshift of 4 also show that injection of a hard-spectrum electron population could explain the broadband flare events (Gokus et al., 2024, 2025). Interestingly, even in the low flux state, the X-ray emission of B3 1239+376 is luminous, which is also seen in several -ray quiet high- blazars (Marcotulli et al., 2020, 2025). Together with the hard X-ray spectra, blazars are believed as powerful MeV beacons in the early cosmos.
The Compton dominance (CD) of B3 1239+376 is indicated as 10, consistent with that found in other high- blazars (e.g., Ghisellini et al., 2011; Ackermann et al., 2017; Gokus et al., 2025). In particular, the CD values of B3 1428+422 and GB 1508+5714 are given as 16 and 23, respectively (Gokus et al., 2024, 2025). CD could be an important probe on the accretion efficiency of the central SMBH, since the dense external photon fields from the accretion system could significantly boost the energy release in the -ray domain (Nalewajko and Gupta, 2017). In the scheme of the blazar sequence (Fossati et al., 1998; Meyer et al., 2011; Prandini and Ghisellini, 2022), more luminous blazars tend to have lower synchrotron peak frequencies and higher CD values. Majority of the hi- blazars, including B3 1239+376, fit this profile. However, note that detection of a high- -ray-emitting BL Lac has been reported, which possesses a relatively high-frequency peaked and a low-Compton dominated SED (Paliya et al., 2020). The construction of an unbiased sample at the high luminosity end of blazars is helpful to critically examine the sequence (Giommi et al., 2012; Keenan et al., 2021).
Meanwhile, an important jet parameter, the bulk Lorentz factor (), is worthy to be mentioned specially. In the classic orientation based AGN unified scheme, the number ratio of the jetted AGNs between the aligned and the apparent populations is suggested as . Thus, is a key parameter to determine the number density of jetted AGNs in the early cosmic time. The value for B3 1239+376 obtained here is 15, in agreement with the results (i. e., 1020) provided for other similar -ray sources (e.g., Ghisellini et al., 2011; Ackermann et al., 2017; Liao et al., 2020; Gokus et al., 2024, 2025). Marcotulli et al. (2020) has carried out an analysis of SEDs of a sample of high-z blazars detected with NuSTAR, though in which non-simultaneous data were adopted, also suggesting 1015. In addition to the jet SED modeling studies, other approaches are used to probe the jet Lorentz factor of high- sources. Searching for the potential evidence of fast variability, especially in rays, can put a severe constraint on the value (e.g., Li et al., 2018). Moreover, detections of the proper motion of the radio jets provide a direct estimation of the essential properties of the jet (e.g., Zhang et al., 2020; Krezinger et al., 2024; Benke et al., 2024). Considering the recent -ray activity of B3 1239+376 revealed by Fermi-LAT, quick follow-up radio interferometry observations that would reveal its jet kinematics are strongly encouraged.
@@
In summary, based on detailed multi-wavelength investigations, B3 1239+376 ( = 3.82) is identified as a -ray emitter, and hence it is concluded as the third most distant -ray-emitting blazar so far, underneath B3 1428+422 ( = 4.72) and GB 1508+5714 ( = 4.3). Analysis of Fermi-LAT data yields a significantly detected (7.7 globally) -ray source cospatial with B3 1239+376 in a half-year period of 2025. It is known as the unique blazar (candidate) within the location uncertainty and an association probability of 0.91 is estimated, which suggest a physical connection between the -ray source and B3 1239+376. Meanwhile, indications of prior -ray activities are also found. Although for the two epochs each alone the statistics are limited, a joint analysis returns a boosted detection significance and the B3 1239+376 remains within the -ray localization error radius. Moreover, in a multi-wavelength perspective, the -ray and infrared fluxes contemporaneously brighten. In particular, emissions in these two windows reach up to the maximal level at the same time, under a total 16-year monitoring. The temporal coincidence helps us to determine the association relationship. Interestingly, accompanying radio core flux increase until Mar. 2018 is observed. In addition to the temporal behaviors, broadband SEDs of B3 1239+376 are plotted and described by a one-zone leptonic radiation scenario. Ejections of fresh emitting electrons into the dissipation region together with an increase of the magnetic field intensity may be responsible for the variations. The soft -ray spectrum, luminous X-ray and -ray emissions, and variability across different bands found in B3 1239+376, are consistent with that found in other similar sources. To further investigate its jet properties, future multi-wavelength campaigns are urged.